OpenStudy (anonymous):

I think its 1 .. 2x + 5 = x – 4 Subtract x from both sides. x + 5 = -4 Minus 5 from both sides x = ____

OpenStudy (littlebird):

-4-5=-9

OpenStudy (anonymous):

oh so it equals -9? lol omg how embressing

OpenStudy (anonymous):

Can you do the whole problem for me so I can understand it ?

OpenStudy (anonymous):

x=-9

OpenStudy (nirmalnema):

2x+5=x-4 2x-x=-5-4 so x=-9

OpenStudy (anonymous):

Oh okay thank you (:

OpenStudy (nirmalnema):

ur welcome